One number is 4 less than 3 times another. If the sum of the numbers is 36, what are the two numbers
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

X+Y=36
X=36-Y
X=3Y-4 THUS WE SUBSTITUTE (3Y-4) FOR X IN THE OTHER EQUATION & WE GET.
(3Y-4)=36-Y
3Y-4=36-Y
3Y+Y=36+4
4Y=40
Y=10 ANSWER NOW SUBSTITUTE 10 FOR Y & SOLVE FOR X
X+10=36
X=36-10
X=26 ANSWER.
PROOF
26=3*10-4
26=30-4
26=26

Answer by stanbon(48558)

One number is 4 less than 3 times another. If the sum of the numbers is 36
Get symbols for the numbers:
Let one be x; let the other by y.
EQUATIONS:
x+y=36
x=3y-4
----------
Solve by substitution:
3y-4+y=36
4y=32
y=8
The x=3*8-4=28
----------
Solution:
x=28; y=8
===============
Cheers,
Stan H.